OpenStudy (anonymous):

If f[x] = Integrate]t/0 4 Sin[x^3] what does it mean to 'give a clean formula' for f'[x] ? latex coming ...

OpenStudy (anonymous):

\[f(x)= \int\limits_{0}^{t} 4 \sin (x^3) dx\]

OpenStudy (astrophysics):

Looks like they want you to use fundamental theorem of calculus part 1

OpenStudy (anonymous):

would that not just be \[\int\limits_{0}^{t} f'(x) = f(t) - f(0)\] and there for f'(x) = 4 sin (x^3) Or is there more to it than that?

OpenStudy (astrophysics):

Kind of all you have to do here is plug in t where x is

OpenStudy (anonymous):

ahhh thats what I did wrong

OpenStudy (anonymous):

thnx

OpenStudy (astrophysics):

So we have \[\int\limits_{0}^{t} 4 \sin(x^3) dx = 4\sin(t^4)=f'(x)\]eas right

OpenStudy (astrophysics):

easy*

OpenStudy (astrophysics):

Np :-)

OpenStudy (anonymous):

had something like d/dx f[t] = f'[x]

OpenStudy (astrophysics):

Sort of haha, basically taking the derivative of integral, which just cancels out the integral, if that makes sense

OpenStudy (astrophysics):

It'll be rare for you to see fundamental theorem of calc part 1 after calc 1 xD

OpenStudy (astrophysics):

But part 2 what you did first is everywhere